### v1

@@ 2.1 Basic setup
Each unit of resource $b\in\mathcal{B}$ is associated with a perishing time $T_{b}\in\mathbb{N}^{+}$ drawn from a known distribution. Items’ perishing times are independent of one another and of the arrival process, and perishing occurs at the end of each round, after items have been allocated to individuals. For $t\in[T]$, we let $P_{t}=\sum_{b\in\mathcal{B}}\mathds{1}\mathopen{}\mathclose{{}\left\{T_{b}=t}\right\}$ denote the number of units of resource perishing in period $t$.
The decision-maker has access to a predicted ordering according to which items perish; we will often refer to this ordering as the allocation schedule. We use $\sigma:\mathcal{B}\rightarrow[B]$ to denote this ordering, i.e., $\sigma(b)$ is rank of $b$ in this ordering. For $b\in[B]$, $\sigma^{-1}(b)$ is used to denote the identity of the $b$th ranked item in $\sigma$, with $\sigma^{-1}(1)$ being the item that comes first in the allocation schedule. While our results allow $\sigma$ to be arbitrary, in Section 5 we investigate natural choices of $\sigma$, such as increasing order of $\mathbb{E}\mathopen{}\mathclose{{}\left[T_{b}}\right]$.
Remark 2.1.
@@
For any time-dependent quantity $Y_{t}$, we define $Y_{\leq t}=\sum_{t^{\prime}\leq t}Y_{t^{\prime}}$, $Y_{\geq t}=\sum_{t^{\prime}\geq t}Y_{t^{\prime}}$, along with their strict analogs. We let $w_{max}=\max_{\theta}w_{\theta}$, $\sigma_{t,\theta}^{2}=\text{Var}(N_{t,\theta})<\infty$, and assume $\rho_{t,\theta}=|N_{t,\theta}-\mathbb{E}[N_{t,\theta}]|<\infty$ almost surely. Finally, let $\mu_{\max}=\max_{t}\mathbb{E}\mathopen{}\mathclose{{}\left[N_{t}}\right],\sigma^{2}_{\min}=\min_{t,\theta}\sigma^{2}_{t,\theta},\sigma^{2}_{\max}=\max_{t,\theta}\sigma^{2}_{t,\theta}$, and $\rho_{\max}=\max_{t,\theta}\rho_{t,\theta}$. We use $\lesssim$ and $\gtrsim$ to denote the fact that inequalities hold up to polynomial factors of $\beta_{avg},|\Theta|,w_{max},\mu_{\max},\sigma^{2}_{\min},\sigma^{2}_{\max}$, $\log T$, and $\log(1/\delta)$. We summarize all notation in Table 5.
@@ 2.2 Notions of fairness and efficiency
The goal is to design a fair and efficient online algorithm that determines the amount to allocate to all $N_{t,\theta}$ individuals in each round $t$, for all $\theta\in\Theta$, given the remaining budget in each round. We assume this amount is allocated uniformly across all $N_{t,\theta}$ individuals of type $\theta$. We use $X_{t,\theta}^{alg}\in\mathbb{R}$ to denote the per-individual amount distributed in period $t$, with $X^{alg}=(X_{t,\theta}^{alg})_{t\in[T]}$.
Our notions of online fairness and efficiency are motivated by the offline notion of Varian Fairness (Varian, , 1974), and are the same as those considered in past works (Sinclair et al., , 2022).
Definition 2.2 (Counterfactual Envy, Hindsight Envy, and Efficiency).
@@
    $\Delta_{\text{\it EF}}\triangleq\max_{t\in[T],\theta\in\Theta}\mathopen{}\mathclose{{}\left|w_{\theta}\mathopen{}\mathclose{{}\left(X_{t,\theta}^{alg}-\frac{B}{N}}\right)}\right|.$ (1)
  • Hindsight Envy:
@@
    $\textsc{Envy}\triangleq\max_{t,t^{\prime}\in[T]^{2},\theta,\theta^{\prime}\in\Theta^{2}}w_{\theta}(X_{t^{\prime},\theta^{\prime}}^{alg}-X_{t,\theta}^{alg}).$ (2)
  • Inefficiency:
@@
    $\Delta_{\text{\it efficiency}}\triangleq B-\sum_{t\in[T],\theta\in\Theta}N_{t,\theta}X_{t,\theta}^{alg}.$ (3)
@@
In the offline setting without perishable goods, Varian, (1974) established that $X^{opt}_{t,\theta}=B/N$ (referred to as the proportional allocation) is the optimal fair and efficient per-individual allocation. Hence, counterfactual envy $\Delta_{EF}$ can be interpreted as a form of regret with respect to this strong no-perishing benchmark, and can be used to characterize the impact of perishability on our algorithm’s performance. Hindsight envy, on the other hand, measures how differently the online algorithm treats any two individuals across time. Finally, the efficiency of the online algorithm, $\Delta_{\text{\it efficiency}}$, measures how wasteful the algorithm was in hindsight. This could happen in two ways: either through spoilage, or because the decision-maker allocated too conservatively throughout the horizon, thus leaving a large number of unspoiled goods unallocated by $T$.
@@
Even in simple settings without perishability, it is known that these metrics are at odds with each other in online settings. To see this, consider the following two scenarios. On the one hand, an algorithm can trivially achieve a hindsight envy of zero by allocating nothing to individuals in any round; this, however, would result in both high counterfactual envy, in addition to maximal inefficiency. On the other hand, a distance to efficiency of zero can trivially be satisfied by exhausting the budget in the first round, at a cost of maximal hindsight envy as individuals arriving at later rounds receive nothing. Sinclair et al., (2022) formalized this tension for the additive utility setting without perishable resources via the following lower bounds.
Theorem 2.3 (Theorems 1 and 2, Sinclair et al., (2022)).
@@
Under any arrival distribution satisfying mild regularity conditions, there exists a problem instance without perishing, such that any algorithm must incur $\Delta_{\text{\it EF}}\gtrsim\frac{1}{\sqrt{T}}$, where $\gtrsim$ drops poly-logarithmic factors of $T$, $\log(1/\delta)$, $o(1)$ terms, and absolute constants. Moreover, any algorithm that achieves $\Delta_{\text{\it EF}}=L_{T}=o(1)$ or $\textsc{Envy}=L_{T}=o(1)$ must also incur waste $\Delta_{\text{\it efficiency}}\gtrsim\min\{\sqrt{T},1/L_{T}\}.$
@@
Since settings without perishable resources are a special case of our setting (e.g., a perishing process with $T_{b}>T$ a.s., for all $b\in\mathcal{B}$), this lower bound holds in our case; the goal then is to design algorithms that achieve this lower bound with high-probability. However, as we will see in the following section, perishing stochasticity is fundamentally distinct from, and more challenging than, demand stochasticity. This difference is particularly salient in regards to the envy-efficiency trade-off.
3 Limits of perishability
@@
In the presence of perishable resources, a decision-maker must contend with two obstacles: $(i)$ the “aggressiveness” of the perishing process, and $(ii)$ errors in the perishing prediction $\sigma$. In this section we formalize the impact of these two challenges. Namely, we identify classes of perishing processes for which there is no hope of achieving the optimal fair allocation, and derive lower bounds on any algorithm’s performance, as a function of the quality of the prediction $\sigma$.
In the remainder of the section, we say that an online algorithm is feasible over a sample path if it does not run out of budget. Note that, if an algorithm is infeasible over a sample path, it necessarily achieves $\Delta_{\text{\it EF}}=\Theta(1)$.
3.1 Restricting the aggressiveness of the perishing process
@@
To see this, consider an instance where all items perish at the end of the first round. There is no hope of achieving low envy in this setting, since there are no items left for arrivals from $t=2$ onwards. The following result establishes that the only perishing time realizations for which $B/N$ is a meaningful benchmark are ones in which the fraction of perished items “lags behind” the proportion of arrivals in each period. We formalize this via the notion of offset-expiry, defined below.
Definition 3.1 (Offset-expiring process).
@@
$\frac{P_{<t}}{B}\leq\frac{N_{<t}}{N}\quad\forall\ t\geq 2.$
Theorem 3.2 establishes that offset-expiry exactly captures the trajectories whereby $B/N$ is a feasible allocation when units are allocated in increasing order of $T_{b}$. We refer to this latter ordering as the hindsight optimal ordering. We defer its proof to Section B.1.
Theorem 3.2.
@@
Thus motivated, for our main algorithmic result we restrict our attention to processes that satisfy the offset-expiry condition with high probability (via a relaxed notion of $\delta$-offset-expiry, see Definition 4.1). We moreover provide verifiable necessary and sufficient conditions for this condition to hold in Section 4.2. From a managerial perspective, the (high-probability) offset-expiry condition provides decision-makers with guidance on the selection of perishable goods to stock at the beginning of the horizon. Within the context of food banks, for instance, it highlights that rejecting perishable goods outright is too severe a policy, and that the reasonable rule of thumb that says “don’t have more spoilage than what you want to give out” is the only correct rule of thumb. It moreover underscores the importance of jointly considering demand and perishing processes in this selection.
Finally, we note that though the remainder of our work is focused on high-probability offset-expiring processes, an interesting future research direction is the question of fairly allocating non-offset-expiring goods under more meaningful notions of envy and inefficiency that don’t penalize for aggressive perishing.
3.2 Unavoidable loss due to errors in allocation schedule
The previous section established that, even under full information about the perishing process, there exist restrictions on the aggressiveness of the process for the optimal fair allocation to be achievable. We next show that, even under offset-expiry, the quality of the allocation schedule $\sigma$ is crucial in determining what any online algorithm can achieve. To see this, consider an instance where $B=T$, $|\Theta|=1$ (with $w_{\theta}=1)$, and $N_{t}=1$ for all $t$. Suppose moreover that exactly one unit perishes in each period (i.e., the perishing process is offset-expiring), but $\sigma$ reverses the true perishing order. In this case, allocating $B/N=1$ to each arrival under $\sigma$ is infeasible, since after $T/2$ rounds the algorithm will have run out of items.

### v2

@@ 2.1 Basic Setup
Each unit of resource $b\in\mathcal{B}$ is associated with a perishing time $T_{b}\in\mathbb{N}^{+}$ drawn from a known distribution. Items’ perishing times are independent of one another and of the arrival process, and perishing occurs at the end of each round, after items have been allocated to individuals. For $t\in[T]$, we let $P_{t}=\sum_{b\in\mathcal{B}}\mathds{1}\left\{T_{b}=t\right\}$ denote the number of units of resource perishing in period $t$.
The decision-maker has access to a predicted ordering according to which items perish; we will often refer to this ordering as the allocation schedule. For example, the decision-maker may order items in increasing order of their expected perishing times, or according to a lower confidence bound on their perishing times. As alluded to in the introduction, within the context of food bank operations, the decision-maker may have access to “best-by” dates specified on food packaging and choose to allocate items with earlier “best-by” dates first. We use $\sigma:\mathcal{B}\rightarrow[B]$ to denote this ordering, i.e., $\sigma(b)$ is rank of $b$ in this ordering. For $b\in[B]$, $\sigma^{-1}(b)$ is used to denote the identity of the $b$th ranked item in $\sigma$, with $\sigma^{-1}(1)$ being the item that comes first in the allocation schedule.
Example 2.1
Consider an instance for which $\mathcal{B}=\{\texttt{a},\texttt{b},\texttt{c}\}$, and let $\sigma$ be such that $\sigma(\texttt{a})=2$, $\sigma(\texttt{b})=3$, and $\sigma(\texttt{c})=1$. Then, $\sigma^{-1}(1)=\texttt{c}$, $\sigma^{-1}(2)=\texttt{a}$, and $\sigma^{-1}(3)=\texttt{b}$.
Remark 2.2
@@
For any time-dependent quantity $Y_{t}$, we define $Y_{\leq t}=\sum_{t^{\prime}\leq t}Y_{t^{\prime}}$, $Y_{\geq t}=\sum_{t^{\prime}\geq t}Y_{t^{\prime}}$, along with their strict analogs. We let $w_{max}=\max_{\theta}w_{\theta}$, $\sigma_{t,\theta}^{2}=\text{Var}(N_{t,\theta})<\infty$, and assume $\rho_{t,\theta}=|N_{t,\theta}-\mathbb{E}[N_{t,\theta}]|<\infty$ almost surely. Finally, let $\mu_{\max}=\max_{t}\mathbb{E}\left[N_{t}\right],\sigma^{2}_{\min}=\min_{t,\theta}\sigma^{2}_{t,\theta},\sigma^{2}_{\max}=\max_{t,\theta}\sigma^{2}_{t,\theta}$, and $\rho_{\max}=\max_{t,\theta}\rho_{t,\theta}$. We use $\lesssim$ and $\gtrsim$ to denote the fact that inequalities hold up to polynomial factors of $\beta_{avg},|\Theta|,w_{max},\mu_{\max},\sigma^{2}_{\min},\sigma^{2}_{\max}$, $\log T$, and $\log(1/\delta)$.
@@ 2.2 Notions of Fairness and Efficiency
The goal is to design a fair and efficient online algorithm that determines the amount to allocate to all $N_{t,\theta}$ individuals in each round $t$, for all $\theta\in\Theta$, given the remaining budget in each round. We assume this amount is allocated uniformly across all $N_{t,\theta}$ individuals of type $\theta$. We use $X_{t,\theta}^{alg}\in\mathbb{R}_{+}$ to denote the per-individual amount distributed in period $t$, with $X^{alg}=(X_{t,\theta}^{alg})_{t\in[T]}$. Finally, we let $B_{t}^{alg}$ denote the remaining budget at the beginning of period $t$, defined inductively as follows:
[…]
where $\textsc{PUA}_{t}^{alg}$ denotes the quantity of unallocated items that perish at the end of round $t$ (i.e., the perished and unallocated items). We moreover let $\mathcal{B}_{t}^{alg}$ be the set of items remaining at the beginning of period $t$.
Our notions of online fairness and efficiency are motivated by the offline notion of Varian Fairness (Varian 1974), and are the same as those considered in past works (Sinclair et al. 2023).
Definition 2.3 (Counterfactual Envy, Hindsight Envy, and Efficiency)
@@
$\bullet\ \text{\emph{Counterfactual Envy}:}$ $\Delta_{\text{\it EF}}$ $\triangleq\max_{t\in[T],\theta\in\Theta}\left|w_{\theta}\left(X_{t,\theta}^{alg}-\frac{B}{N}\right)\right|$ (2)
@@
$\bullet\ \text{\emph{Hindsight Envy}:}$ Envy $\triangleq\max_{t,t^{\prime}\in[T]^{2},\theta,\theta^{\prime}\in\Theta^{2}}w_{\theta}(X_{t^{\prime},\theta^{\prime}}^{alg}-X_{t,\theta}^{alg})$ (3)
@@
$\bullet\ \text{\emph{Inefficiency}:}$ $\Delta_{\text{\it efficiency}}$ $\triangleq B-\sum_{t\in[T],\theta\in\Theta}N_{t,\theta}X_{t,\theta}^{alg}$ (4)
@@
In the offline setting without perishable goods, Varian (1974) established that $X^{opt}_{t,\theta}=B/N$ (referred to as the proportional allocation) is the optimal fair and efficient per-individual allocation. Hence, counterfactual envy $\Delta_{EF}$ can be interpreted as a form of regret with respect to this strong no-perishing benchmark, and can be used to characterize the impact of perishability on our algorithm’s performance. Hindsight envy, on the other hand, measures how differently the online algorithm treats any two individuals across time. Finally, the efficiency of the online algorithm, $\Delta_{\text{\it efficiency}}$, measures how wasteful the algorithm was in hindsight. This could happen in two ways: either through spoilage, or because the decision-maker allocated too conservatively throughout the horizon, thus leaving a large number of unspoiled goods unallocated by $T$.
@@
Even in simple settings without perishability, it is known that these metrics are at odds with each other in online settings. To see this, consider the following two scenarios. On the one hand, an algorithm can trivially achieve a hindsight envy of zero by allocating nothing to individuals in any round; this, however, would result in both high counterfactual envy, in addition to maximal inefficiency. On the other hand, a distance to efficiency of zero can trivially be satisfied by exhausting the budget in the first round, at a cost of maximal hindsight envy as individuals arriving at later rounds receive nothing. Sinclair et al. (2023) formalized this tension for the additive utility setting without perishable resources via the following lower bounds.
Theorem 2.4 (Theorems 1 and 2, Sinclair et al. (2023))
@@
Under any arrival distribution satisfying mild regularity conditions, there exists a problem instance without perishing, such that any algorithm must incur $\Delta_{\text{\it EF}}\gtrsim\frac{1}{\sqrt{T}}$. Moreover, any algorithm that achieves $\Delta_{\text{\it EF}}=L_{T}=o(1)$ or $\textsc{Envy}=L_{T}=o(1)$ must also incur $\Delta_{\text{\it efficiency}}\gtrsim\min\{\sqrt{T},1/L_{T}\}.$
@@
Since settings without perishable resources are a special case of our setting (e.g., a perishing process with $T_{b}>T$ a.s., for all $b\in\mathcal{B}$), this lower bound holds in our case; the goal then is to design algorithms that achieve this lower bound with high-probability. However, as we will see in Section 3, perishing stochasticity is fundamentally distinct from, and more challenging than, demand stochasticity. This difference is particularly salient in regards to the envy-efficiency trade-off.
2.3 Discussion of Modeling Assumptions
We conclude the section with a discussion of our modeling assumptions.
Age-dependent quality considerations.
While our model does not explicitly account for freshness considerations (i.e., allocating an item that perishes tomorrow is considered to be the same “good” outcome as allocating an item that perishes in 10 days), one natural way to take such considerations into account would be to define a pseudo-perishing time $\widetilde{T}_{b}=T_{b}-a_{b}$, where $a_{b}>0$ is an item-dependent constant that guarantees a minimum amount of time that an item lasts in the household once it is allocated. (In practice, the decision-maker may want to set $a_{b}$ in a category-specific way, such that the allocated item meets a minimum quality threshold at the time of allocation. For instance, milk can last 5-7 days after its “sell-by” date if stored properly in a refrigerator (U.S. Dairy 2025). In this case, it may be desirable to set $a_{b}=3$ in order to ensure that a household can consume the milk for at least 3 days.) Defining all metrics with respect to these pseudo-perishing times, our results would continue to hold, thereby guaranteeing some notion of freshness in our end allocation. [^1] If $\sigma$ is chosen according to the distribution of the pseudo-perishing time $\widetilde{T}_{b}$, the choice of $a_{b}$ may also have an impact on $\sigma$. For instance, if items are allocated in increasing order of $\mathbb{E}[\widetilde{T}_{b}]$, items for which $a_{b}$ is larger (i.e., those for which there is a quality premium) will be allocated earlier. If the decision-maker allocates items according to the variance of $\widetilde{T}_{b}$, on the other hand, given that $a_{b}$ is a constant, this choice would have no impact on $\sigma$.
Alternatively, one could encode freshness considerations by modeling time-dependent utilities for each item. While this would be an interesting direction to explore theoretically, it is less practically relevant with respect to our motivating application of food banks, given that the quantity of goods allocated — as opposed to an estimate of utility-based social welfare —- remains the most important metric reported by these organizations (Prendergast 2022, Orgut et al. 2016).
On the allocation schedule $\sigma$.
In this work we assume the allocation schedule $\sigma$ is an input to the model, as opposed to a lever available to the decision-maker. While treating $\sigma$ as a decision variable would improve both the envy and inefficiency of our algorithm, the reason for this decoupling is two-fold. From a practical perspective, one can imagine that decision-makers have a preference for using interpretable allocation schedules, such as according to expected perishing time or “best-by” date. In contrast, having an allocation schedule depend on the specifics of the guardrail algorithm would be more challenging to communicate to stakeholders. This would be problematic in instances where spoilage occurred, and a unit with a later “best-by” date or expected perishing time was allocated before the unit that spoiled. Such bad events are harder to justify when $\sigma$ is an output of the decision-maker’s optimization problem. By allowing $\sigma$ to be an input to the model, our framework can instead be used by the decision-maker in practice to test the performance of various interpretable allocation schedules, and heuristically determine the one that yields the strongest performance, as we do in our numerical experiments.
We moreover posit that treating $\sigma$ as a decision variable would be analytically challenging, as well as a computationally intractable task. Namely, as we will see in Section 3, optimizing over $\sigma$ reduces to an optimization over $N!$ orderings, which cannot be done efficiently. We however make some progress toward characterizing “high-quality” allocation schedules by $(i)$ providing sufficient conditions on the relationship between $\sigma$ and the perishing distribution for our algorithm’s spoilage to vanish asymptotically (Proposition 4.19), and $(ii)$ numerically showing that allocating units in increasing order of a lower confidence bound on their perishing time is a robust choice for $\sigma$ (Section 5.4). We leave the derivation of structural results surrounding the optimal allocation schedule (which we conjecture will, in general, be highly distribution-dependent), as well as the design of approximately optimal allocation schedules, as interesting open questions.
Finally, we note that the ranking of units is a sufficient statistic for the model we consider in this paper (as opposed to, say, granular forecasts of their perishing times). To see this, note that given a sequence of allocation quantities, the only question that remains for the decision-maker is which units she should allocate in each period. This reduces to the task of finding a ranking of units that minimizes spoilage, given the allocation quantities. One can easily see this when perishing times are known and deterministic; in this case, given an allocation quantity in each period, knowledge of the actual perishing times provides no additional advantage over the ranking that is induced by these true perishing times. This similarly holds for any forecast of units’ perishing times treated as ground truth.
3 Fundamental Limits of Perishability: Lower Bounds
@@
In the presence of perishable resources, a decision-maker must contend with two obstacles: $(i)$ the aggressiveness of the perishing process, and $(ii)$ errors in the perishing prediction $\sigma$. In this section we formalize the impact of these two challenges. Namely, we identify classes of perishing processes for which there is no hope of achieving the optimal fair allocation, $B/N$, and derive lower bounds on any algorithm’s performance as a function of a salient quantity that captures both the aggressiveness of perishing, as well as the quality of the prediction $\sigma$. In the remainder of the section, we say that an online algorithm is feasible over a sample path if it does not run out of budget. Note that, if an algorithm is infeasible over a sample path, it necessarily achieves $\Delta_{\text{\it EF}}=\Theta(1)$.
3.1 On the Aggressiveness of the Perishing Process
@@
To see this, consider an instance where all items perish at the end of the first round. There is no hope of achieving low envy in this setting, since there are no items left for arrivals from $t=2$ onwards. The following result establishes that the only perishing time realizations for which $B/N$ is achievable are ones in which the fraction of perished items “lags behind” the proportion of arrivals in each period. We formalize this via the notion of offset-expiry, defined below.
Definition 3.1 (Offset-expiring process)
@@
$\frac{P_{<t}}{B}\leq\frac{N_{<t}}{N}\quad\forall\ t\geq 2.$ (5)
We provide an example below.
Example 3.2 (Continued from Example 2.1)
Consider a setting where $T=B=3$, with $N_{1}=1$, $N_{2}=2$, and $N_{3}=1$ (and $N=4$ as a result). Suppose moreover $|\Theta|=1$, with $w_{\theta}=1$. For this sample path, we have:
$\frac{N_{<t}}{N}=\begin{cases}\frac{1}{4}\quad&\text{for }t=2\\ \frac{3}{4}&\text{for }t=3.\end{cases}$ (6)
Applying Equation 5 to this sample path, a process is offset-expiring if and only if:
$P_{<t}\leq\begin{cases}0\quad&\text{for }t=2\\ 2&\text{for }t=3.\end{cases}$ (7)
In words, for any offset-expiring process in this example, we must have $T_{b}\geq 2$ almost surely for all $b\in\mathcal{B}$, with no more than two units perishing at $t=2$ in the worst case.
Theorem 3.3 establishes that offset-expiry exactly captures the trajectories whereby $B/N$ is a feasible allocation when units are allocated in increasing order of $T_{b}$. We refer to this latter ordering as the hindsight optimal ordering. We defer its proof to Appendix 7.
Theorem 3.3
@@
From a managerial perspective, the offset-expiry condition provides decision-makers with guidance on the selection of perishable goods to stock at the beginning of the horizon. Within the context of food banks, for instance, it highlights that rejecting perishable donations outright (as some food banks do) is too severe a policy; rather, the reasonable rule of thumb that says “don’t accept food if it will spoil faster than the rate at which you allocate it” not only suffices, but is the only correct rule of thumb. It moreover underscores the importance of jointly considering the demand and perishing distributions in this selection. We provide necessary and sufficient conditions for offset-expiry to hold with high probability in Section 4.4.
3.2 Unavoidable Loss Due to Perishing
The previous section established that, even under full information about the perishing process, there exist aggressive perishing distributions for which the proportional allocation $B/N$ is not achievable. The following example illustrates how both the inherent aggressiveness of the exogenous perishing process, as well as the quality of the allocation schedule $\sigma$, play an important role in determining what an algorithm can achieve.
Example 3.4 (Continued from Examples 2.1 and 3.2)
Consider the following perishing distributions:
$T_{\texttt{a}}=\begin{cases}1\quad\text{w.p. }1/2\\ 3\quad\text{w.p. }1/2\end{cases}\quad T_{\texttt{b}}=\begin{cases}2\quad\text{w.p. }1/2\\ 4\quad\text{w.p. }1/2\end{cases}\quad T_{\texttt{c}}=\begin{cases}1\quad\text{w.p. }1/2\\ 2\quad\text{w.p. }1/2\end{cases}.$ (8)
Recall, $N_{1}=1$, $N_{2}=2$, and $N_{3}=1$; moreover, $\sigma(\texttt{a})=2$, $\sigma(\texttt{b})=3$, and $\sigma(\texttt{c})=1$. Note that $\sigma$ orders the units in increasing order of expected perishing time for this instance.
By Example 3.2, this process is not offset-expiring for any sample path in which $T_{\texttt{a}}=1$ or $T_{\texttt{c}}=1$. By Theorem 3.3, then, the proportional allocation of $B/N=3/4$ is only feasible on sample paths for which $T_{\texttt{a}}=3$ and $T_{\texttt{c}}=2$, under the hindsight optimal ordering. Even over such sample paths, however, the possibility remains that $B/N$ is infeasible under $\sigma$. To see this, consider the case where $T_{\texttt{a}}=3$, $T_{\texttt{c}}=2$, and $T_{\texttt{b}}=2$. By allocating $3/4$ in each period according to ordering $\sigma$, $2.25$ units will have been allocated by the end of period $2$ (i.e., the entirety of $\texttt{c},\texttt{a}$, and $0.25$ units of b). However, the remainder of b perishes at the end of period $2$, leaving nothing for the individual who arrives in period $3$.